# How Aligned are Human Chart Takeaways and LLM Predictions? A Case Study on Bar Charts with Varying Layouts

### v1

@@ Abstract
Large Language Models (LLMs) have been adopted for a variety of visualizations tasks, but how far are we from perceptually aware LLMs that can predict human takeaways? Graphical perception literature has shown that human chart takeaways are sensitive to visualization design choices, such as spatial layouts. In this work, we examine the extent to which LLMs exhibit such sensitivity when generating takeaways, using bar charts with varying spatial layouts as a case study. We conducted three experiments and tested four common bar chart layouts: vertically juxtaposed, horizontally juxtaposed, overlaid, and stacked. In Experiment 1, we identified the optimal configurations to generate meaningful chart takeaways by testing four LLMs, two temperature settings, nine chart specifications, and two prompting strategies. We found that even state-of-the-art LLMs struggled to generate semantically diverse and factually accurate takeaways. In Experiment 2, we used the optimal configurations to generate 30 chart takeaways each for eight visualizations across four layouts and two datasets in both zero-shot and one-shot settings. Compared to human takeaways, we found that the takeaways LLMs generated often did not match the types of comparisons made by humans. In Experiment 3, we examined the effect of chart context and data on LLM takeaways. We found that LLMs, unlike humans, exhibited variation in takeaway comparison types for different bar charts using the same bar layout. Overall, our case study evaluates the ability of LLMs to emulate human interpretations of data and points to challenges and opportunities in using LLMs to predict human chart takeaways.
@@ Introduction
Large language models (LLMs) have recently taken the world by storm due to their remarkable ability to generate coherent text and follow instructions. Trained on vast text corpora, LLMs have the potential to encode rich visualization knowledge. Thus, they offer exciting new possibilities for visualization tasks. For instance, Chen et al. [11] utilized Codex [10] to transform natural language queries into structurally parameterized SQL, which was used to generate interactive visualization interfaces. More powerful LLMs further simplify visualization generation. For example, GPT-4 allows users to specify chart requirements and directly produces high-quality visualizations in many instances [9].
So can LLMs effectively leverage its knowledge base to account for visualization design affordances? In this paper, we compare the sensitivity of LLMs to changes in visualization design to that of humans’ when generating visualization takeaways. Specifically, we explore whether LLMs are perceptually aware of design manipulations like humans are by examining bar charts as a case study. Existing work by Xiong et al. [39] has examined in detail how common layouts affect the type of comparisons people make when reading bar charts. The researchers generated two-by-three bar charts using two datasets and visualized them in four layouts (vertically juxtaposed, horizontally juxtaposed, overlaid, and stacked, as shown in Figure 1). They collected a corpus of natural language takeaways via crowdsourcing, classifying the types of comparisons human readers made upon seeing these bar charts into 12 categories. Depending on the spatial layout of the bar chart, a human reader can be more or less likely to make a comparison of a specific type. For example, for a dataset depicting the revenue of three stores (1, 2, and 3) in two companies (A and B), people were more likely to compare the revenue of Company A’s Store 1 to that of Company B’s Store 1 when reading an adjacent bar chart, as shown in Figure 3, but they do not make that comparison when they read an overlaid or stacked version of the bar chart.
@@
Experiment 1: Identifying the optimal configurations. We experimented with different configurations for generating chart takeaways across four LLMs, two temperatures, nine chart specifications, and two prompting strategies in zero-shot and one-shot settings. We identified representing charts with ggplot2, utilizing GPT-4 with a temperature of 0, and employing the guided discovery strategy as the optimal configuration for the zero-shot setting and representing charts with Vega-Lite and image, using GPT-4V with a temperature of 0.7, and employing the baseline strategy as the optimal configuration for the one-shot setting. In addition, we found that state-of-the-art LLMs could struggle to generate semantically diverse and factually accurate takeaways.
Experiment 2: Human-alignment of LLM-generated takeaways. Using the optimal configurations identified in Experiment 1, we generated 30 chart takeaways across four bar chart layouts using two datasets. We found that LLMs, unlike humans, were generally insensitive to bar layouts, while providing an in-context example helped align model generations in some cases.
Experiment 3: Effect of context and data. If LLMs are not perceptually sensitive to bar layouts, what are they sensitive to? We observed that the comparison types in LLM takeaways varied greatly between charts with the same layout but different contexts and data values. We showed that humans did not exhibit this sensitivity, thus revealing inconsistency as another weakness of LLMs. We further demonstrated that context affected comparison types more than data did for most layouts.
Our work is a first step towards understanding the perceptual awareness of LLMs for visualization tasks. In our case study on bar charts, we not only discovered that LLMs were generally incapable of writing human-aligned takeaways, but they sometimes also generated semantically repetitive and factually incorrect takeaways. We hope our work can motivate investigation into other visualization types and additional dimensions of perceptual awareness. By exposing the weaknesses of LLMs in reading charts, our work outlines challenges in employing LLMs for visualization tasks and informs future LLM development.
@@ 2.3.2 Guided Discovery
The guided discovery strategy was proposed by Ko et al. [21] in their framework generating natural language datasets from Vega-Lite specifications and drew from chain-of-thought prompting [35] and educational psychology [7]. In guided discovery, the user provides scaffolding and poses key questions in the prompt to help the model reason and extract insights. We instructed the model to reason step-by-step: first, we provided the chart and asked the model to extract the chart type and the variables depicted for grounding purposes (scaffolding); then, we instructed the model to attend to bar heights and reason about the pattern shown; finally, we requested 30 semantically distinct takeaways from the visualization. See the following zero-shot example:
@@ 2.3.3 Prompts for one-shot settings
The examples in the previous sections are templates for zero-shot settings. For one-shot settings, we provided a sample bar chart using the same spatial layout and specified in the same format as the test case and approximately 150 human takeaways before asking the model for takeaways on the test case. We also tested both the baseline and the guided discovery strategies for one-shot prompts. We largely adhered to the task descriptions and system prompts previously identified. Please refer to the supplemental materials for our one-shot templates.
@@ 2.4 Experiment 1: Procedure and Setup
The main goal of Experiment 1 is to identify optimal configurations of the four parameters outlined in Section 2 for zero-shot and one-shot settings. To this end, we prompted LLMs to generate 30 semantically diverse takeaways for each input chart specification. As discussed in the previous sections, we evaluated four LLMs, two temperatures, nine chart representations [^5] Note that we tested chart specifications involving bitmap images only on GPT-4V since it was the only LLM with vision capability among the four tested., and two prompting strategies on two datasets in both zero-shot and one-shot settings. Given the large number of configurations, we only experimented on the adjacent layout in Experiment 1 to derive optimal configurations. This resulted in a total of 136 trials across both zero-shot and one-shot settings. To ensure the integrity and independence of our results, we conducted each trial through separate API calls. Finally, we evaluated the takeaways according to the procedure in Section 2.5 and identified the optimal configurations.
@@ 2.5 Experiment 1: Evaluation Approaches
Figure 4: Distribution of the average cluster count for each configuration (broken down by the LLM type) for the zero-shot and one-shot settings. We reviewed the accuracy of the top 25% (see Section 2.5.1), corresponding to a threshold of 21 for the zero-shot setting and a threshold of 19.5 for the one-shot setting.
@@ 2.5.1 Semantic Diversity
Semantic diversity measures the semantic variety of takeaways, via a “cluster count” proxy. We passed the takeaways for each configuration to GPT-4 and instructed it to cluster them based on their semantics. Take Figure 3 as an example, the takeaway “the revenue of Company A at Store 3 is less than that of Company B” is semantically equivalent to “the revenue of Company B at Store 3 is more than that of Company A”. Therefore, we consider these takeaways to be in the same cluster. These takeaways have different meanings from “Store 2 generates the highest revenue for Company A”, and thus we consider them to be from different clusters. Hence, for a set of takeaways, the higher the cluster count, the more semantically diverse it is.
Since the prompts request 30 semantically distinct takeaways, we first filtered out configurations that generated semantically repetitive takeaways. Recognizing that in-context examples could influence the generated outcomes, we conducted the filtering process separately for zero-shot and one-shot settings. To verify that the cluster counts returned by GPT-4 were reasonable measures of semantic diversity, we selected 10 configurations at the $10^{th}$, $20^{th}$, , and $100^{th}$ percentile of cluster count, manually clustered them, and compared machine and human-generated cluster counts.
@@ 2.5.2 Factual Accuracy
A set of takeaways is considered more factually accurate than another set if it has a higher percentage of correct takeaways. When the percentages of correct takeaways are equal, the set with the lower percentage of factually inaccurate takeaways is deemed more accurate.
@@ 2.6.1 Confirming GPT-4’s Clustering Results
We first verified that the cluster counts generated by GPT-4 were reasonably accurate. After sampling ten sets of takeaways with a diverse range of GPT-4-generated cluster counts following the procedure outlined in Section 2.5.1 and manually clustering them, we performed linear regression to assess the relationship between the GPT-4-generated cluster counts and the manually obtained cluster counts. Results indicated that GPT-4 cluster counts indeed predicted manual cluster counts quite well ($r^{2}=0.83$, $p<0.001$). Hence, we concluded that the cluster counts generated by GPT-4 were good proxies for semantic diversity.
@@ 2.6.2 Semantic Diversity by Configuration
After retaining the top 25% of configurations as described in Section 2.5.1, we were left with configurations with an average cluster count on the two sets of takeaways of above 21 in the zero-shot setting and those with an average above 19.5 in the one-shot setting. This procedure left us with 19 configurations to code for accuracy in the zero-shot setting and 18 in the one-shot setting. In both settings, Gemini 1.0 Pro produced the highest number of semantically diverse takeaways.
@@ 2.6.3 Accuracy by Configuration
Table 2.6.3 and Table 2.6.3 show the configurations in the top five in terms of average factual accuracy for the zero-shot and one-shot settings, respectively. While there were a mix of temperatures, chart representations, and prompting strategies in the top five, all configurations were notably either generated by GPT-4 or GPT-4V, which suggests the GPT-4 family’s dominance over other LLMs in generating accurate takeaways. We also found that Gemini 1.0 Pro, while capable of writing semantically diverse takeaways, struggled in factual accuracy. In fact, most configurations involving Gemini 1.0 Pro averaged below 33.3% in accuracy and the majority of takeaways were ambiguous. Also underperforming in accuracy was GPT-3.5, which often produced more factually inaccurate takeaways than accurate ones. Overall, our results show that generating accurate chart takeaways is still a challenging task for LLMs.
@@ 3 Experiment 2: Alignment of Human & LLM Takeaways
In Experiment 1, we generated takeaways from adjacent bar charts and identified optimal configurations for zero-shot and one-shot settings. Next, we started to more generally understand whether LLMs are influenced by the spatial layouts of bars like humans are. We tested all four layouts in Figure 1, using the optimal configurations identified in Experiment 1. We prompted LLMs to generate 30 takeaways for each visualization as if it were 30 people in both zero-shot and one-shot settings. In addition to coding takeaways for accuracy, we classified them along the taxonomy Xiong et al.developed, which captures 12 types of comparison takeaways can make (denoted C1, C2, , C12, see Section 2.2). By comparing the distributions of LLM-generated takeaways vs. human-written takeaways along these comparison types, we can understand to what extent takeaways from LLMs emulate those from humans.
@@ 3.3 Experiment 2: Results
Figure 5: Distributions of human, LLM zero-shot, and LLM one-shot comparison types for each layout. LLM one-shot distributions are generally closer to human ones than zero-shot distributions.
@@ 3.3.2 Alignment of Comparison Types
Table 3.3.2 further details the top three most frequent comparison types for each layout from each source while Table 3.3.2 shows precision@3. Compared to zero-shot takeaways, one-shot takeaways attained higher precision@3 for two layouts and equal precision@3 for the remaining two, indicating that the presence of example takeaways tended to allow LLMs to better capture the most afforded comparison types. While LLMs were generally good at predicting frequent comparison types across all layouts, such as C1, C11, and C12, when rarer comparison types made it to the top three, such as C3 and C4, LLMs failed to predict them. However, the frequent appearance of a less common comparison type in other layouts within the top ranks for a specific layout suggests a unique affordance by that layout for such comparisons. Failure by the LLM to capture these affordances is further evidence that LLM comparison types are generally insensitive to bar layout.
@@ 4.1 Experiment 3: Procedure, Setup, and Evaluation
To answer the second question, we generated eight new bar charts by modifying the original eight stimuli from Xiong et al.For each chart, we preserved its context while introducing a new dataset. For example, if the existing two stacked bar charts visualized “context 1 and dataset 1” and “context 2 and dataset 2”, we created two new stacked bar charts visualizing “context 1 and dataset 2” and “context 2 and dataset 1”. Next, we generated zero-shot takeaways using the optimal configuration established in Experiment 1, and the same coder from Experiment 2 coded these takeaways for comparison types. We then visualized the LLM comparison type distributions for the four example charts for each layout. To examine the effect of contexts on distributions, we computed the TVD’s between the distributions for each pair of charts showing the same dataset but different contexts and calculated the average. To examine the effect of data on distributions, we computed the TVD’s between the distributions for each pair of charts showing the same context but different datasets and calculated the average.

### v2

@@ Abstract
Large Language Models (LLMs) have been adopted for a variety of visualizations tasks, but how far are we from perceptually aware LLMs that can predict human takeaways? Graphical perception literature has shown that human chart takeaways are sensitive to visualization design choices, such as spatial layouts. In this work, we examine the extent to which LLMs exhibit such sensitivity when generating takeaways, using bar charts with varying spatial layouts as a case study. We conducted three experiments and tested four common bar chart layouts: vertically juxtaposed, horizontally juxtaposed, overlaid, and stacked. In Experiment 1, we identified the optimal configurations to generate meaningful chart takeaways by testing four LLMs, two temperature settings, nine chart specifications, and two prompting strategies. We found that even state-of-the-art LLMs struggle to generate semantically diverse and factually accurate takeaways. In Experiment 2, we used the optimal configurations to generate 30 chart takeaways each for eight visualizations across four layouts and two datasets in both zero-shot and one-shot settings. Compared to human takeaways, we found that the takeaways LLMs generate often do not match the types of comparisons made by humans. In Experiment 3, we examined the effect of chart context and data on LLM takeaways. We found that LLMs, unlike humans, exhibit variation in takeaway comparison types for different bar charts using the same bar layout. Overall, our case study evaluates the ability of LLMs to emulate human interpretations of data and points to challenges and opportunities in using LLMs to predict human chart takeaways.
@@ Introduction
Large language models (LLMs) have recently taken the world by storm due to their remarkable ability to generate coherent text and follow instructions. Trained on vast text corpora, LLMs have the potential to encode rich visualization knowledge. Thus, they offer exciting new possibilities for visualization tasks. For instance, Chen et al. [11] utilized Codex [10] to transform natural language queries into structurally parameterized SQL, which is used to generate interactive visualization surfaces. More powerful LLMs further simplify visualization generation. For example, GPT-4 allows users to specify chart requirements and directly produces high-quality visualizations in many instances [9].
So can LLM effectively leverage its knowledge base to account for visualization design affordances? In this paper, we evaluate the sensitivity of LLMs to changes in visualization design compared to humans, when generating visualization takeaways. Specifically, we explore whether LLMs are perceptually aware of design manipulations like humans are by examining bar charts as a case study. Existing work by Xiong et al. [39] has examined in detail how common layouts affect the type of comparisons people make when reading bar charts. The researchers generated two-by-three bar charts using two datasets and visualized them in four layouts (vertically juxtaposed, horizontally juxtaposed, overlaid, and stacked, as shown in Figure 1). They collected a corpus of natural language takeaways via crowdsourcing, classifying the types of comparisons human readers made upon seeing these bar charts into 12 categories. Depending on the spatial layout of the bar chart, a human reader can be more or less likely to make a comparison of a specific type. For example, for a dataset depicting the revenue of three stores (1, 2, and 3) in two companies (A and B), people were more likely to compare the revenue of Company A’s Store 1 to that of Company B’s Store 1 when reading an adjacent bar chart, as shown in Figure 3, but they do not make that comparison when they read an overlaid or stacked version of the bar chart.
@@
Experiment 1: Identifying the optimal configurations. We experimented with different configurations for generating chart takeaways across four LLMs, two temperatures, nine chart specifications, and two prompting strategies in zero-shot and one-shot settings. We identified representing charts with ggplot2, utilizing GPT-4 with a temperature of 0, and employing the guided discovery strategy as the optimal configuration for the zero-shot setting and representing charts with Vega-Lite and image, using GPT-4V with a temperature of 0.7, and employing the baseline strategy as the optimal configuration for the one-shot setting. In addition, we found that state-of-the-art LLMs can struggle to generate semantically diverse and factually accurate takeaways.
Experiment 2: Hhuman-LLM-alignment of generated takeaways. Using the optimal configurations identified in Experiment 1, we generated 30 chart takeaways across four bar chart layouts using two datasets. We found that LLMs, unlike humans, are generally insensitive to bar layouts, while providing an in-context example helps align model generations in some cases.
Experiment 3: Effect of context and data. If LLMs are not perceptually sensitive to bar layouts, what are they sensitive to? We observed that the comparison types in LLM takeaways vary greatly between charts with the same layout but different contexts and data values. We showed that humans do not exhibit this sensitivity, thus revealing inconsistency as another weakness of LLMs. We further demonstrated that context affects comparison types more than data does for most layouts.
Our work is a first step towards understanding the perceptual awareness of LLMs for visualization tasks. In our case study on bar charts, we not only discovered that LLMs are generally incapable of writing human-aligned takeaways, but they sometimes also generate semantically repetitive and factually incorrect takeaways. We hope our work can motivate investigation into other visualization types and additional dimensions of perceptual awareness. By exposing the weaknesses of LLMs in reading charts, our work outlines challenges in employing LLMs for visualization tasks and informs future LLM development.
@@ 2.3.2 Guided Discovery
The guided discovery strategy was proposed by Ko et al. [21] in their framework generating natural language datasets from Vega-Lite specifications and drew from chain-of-thought prompting [35] and educational psychology [7]. In guided discovery, the user provides scaffolding and poses key questions in the prompt to help the model reason and extract insights. We instructed the model to reason step-by-step: first, we asked the model to extract the chart type and the variables depicted for grounding purposes (scaffolding); then, we instructed the model to attend to bar heights and reason about the pattern shown; finally, we requested 30 semantically distinct takeaways from the visualization and included the chart representation. See the following zero-shot example:
@@ 2.3.3 Prompts for one-shot settings
The examples in the previous sections are instantiations of templates for zero-shot settings. For one-shot settings, we provided a sample bar chart using the same spatial layout and specified in the same format as the test case and approximately 150 human takeaways before asking the model for takeaways on the test case. We also tested both the baseline and the guided discovery strategies for one-shot prompts. We largely adhered to the task descriptions and system prompts previously identified. Please refer to the supplemental materials for our one-shot templates.
@@ 2.4 Experiment 1: Procedure and Setup
The main goal of Experiment 1 is to identify optimal configurations of the four parameters outlined in Section 2 for zero-shot and one-shot settings. To this end, we prompted LLMs to generate 30 semantically diverse takeaways for each input chart specification. As discussed in the previous sections, we evaluated four LLMs, two temperatures, nine chart representations [^5] Note that we tested chart specifications involving bitmap images only on GPT-4V since it was the only LLM with vision capability among the four tested., and two prompting strategies on two datasets in both zero-shot and one-shot settings. Given the large number of configurations, we only experimented on the adjacent layout in Experiment 1 to derive optimal configurations. This resulted in a total of 272 trials across both zero-shot and one-shot settings. To ensure the integrity and independence of our results, we conducted each trial through separate API calls. Finally, we evaluated the takeaways according to the procedure in Section 2.5 and identified the optimal configurations.
@@ 2.5 Experiment 1: Evaluation Approaches
Figure 4: Distribution of the average cluster count for each configuration (broken down by the LLM type) for the zero-shot and one-shot setting. We reviewed the accuracy of the top 25% (see Section 2.5.1), corresponding to thresholds of 21 and 19.5 for the zero-shot and one-shot settings.
@@ 2.5.1 Semantic Diversity
Semantic diversity measures the semantic variety of takeaways, via a “cluster count” proxy. We passed the takeaways for each configuration to GPT-4 and instructed it to cluster them based on their semantics. Take Figure 3 as an example, the takeaway “the revenue of Company A at Store 3 is less than that of Company B” is semantically equivalent to “the revenue of Company B at Store 3 is more than that of Company A”. Therefore, we consider these takeaways to be in the same cluster. These takeaways mean different things from “Store 2 generates the highest revenue for Company A”, and thus we consider them to be from different clusters. Hence, for a set of takeaways, the higher the cluster count, the more semantically diverse it is.
Since the prompts request 30 semantically distinct takeaways, we first filtered out configurations that generated semantically repetitive takeaways. Recognizing that in-context examples could influence the generated outcomes, we conducted the filtering process separately for zero-shot and one-shot settings. To verify that the cluster counts returned by GPT-4 are reasonable measures of semantic diversity, we selected 10 configurations at the $10^{th}$, $20^{th}$, , and $100^{th}$ percentile of cluster count, manually clustered them, and compared machine and human-generated cluster counts.
@@ 2.5.2 Factual Accuracy
A set of takeaways is considered more factually accurate if it has a higher percentage of correct takeaways compared to another set. When the percentage of correct takeaways is equal, the set with the lower percentage of factually inaccurate takeaways is deemed more accurate.
@@ 2.6.1 Confirming GPT-4’s Clustering Results
We first verified that the cluster counts generated by GPT-4 are reasonably accurate. After sampling ten sets of takeaways with a diverse range of GPT-4-generated cluster counts following the procedure outlined in Section 2.5.1 and manually clustering them, we performed linear regression to assess the relationship between the GPT-4-generated cluster counts and the manually obtained cluster counts. Results indicated that GPT-4 cluster counts indeed predict manual cluster counts quite well ($r^{2}=0.83$, $p<0.001$). Hence, we concluded that the cluster counts generated by GPT-4 are good proxies for semantic diversity.
@@ 2.6.2 Semantic Diversity by Configuration
After retaining the top 25% of configurations as described in Section 2.5.1, we are left with configurations with an average cluster count on the two sets of takeaways of above 21 in the zero-shot setting and those with an average above 19.5 in the one-shot setting. This procedure left us with 19 configurations to code for accuracy in the zero-shot setting and 18 in the one-shot setting. In both settings, Gemini 1.0 Pro produces the highest number of semantically diverse takeaways.
@@ 2.6.3 Accuracy by Configuration
Table 2.6.3 and Table 2.6.3 show the configurations in the top five in terms of average factual accuracy for the zero-shot and one-shot settings, respectively. While there are a mix of temperatures, chart representation, and prompting strategies in the top five, all configurations are notably either generated by GPT-4 or GPT-4V, which suggests the GPT-4 family’s dominance over other LLMs in generating accurate takeaways. We also find that Gemini 1.0 Pro, while capable of writing semantically diverse takeaways, struggles in factual accuracy. In fact, most configurations involving Gemini 1.0 Pro average below 33.3% in accuracy and the majority of takeaways are ambiguous. Also underperforming in accuracy is GPT-3.5, which often produces more factually inaccurate takeaways than accurate ones. Overall, our results show that generating accurate chart takeaways is still a challenging task for LLMs.
@@ 3 Experiment 2: Alignment of Human & LLM Takeaways
In Experiment 1, we generated takeaways from adjacent bar charts and identified optimal configurations for the zero-shot and one-shot settings. Next, we started to more generally understand whether LLMs are influenced by the spatial layouts of bars like humans are. We tested all four layouts in Figure 1, using the optimal configurations identified in Experiment 1. We prompted LLMs to generate 30 takeaways for each visualization as if it were 30 people in both zero-shot and one-shot settings. In addition to coding takeaways for accuracy, we classified them along the taxonomy Xiong et al.developed, which captures 12 types of comparison each takeaway makes (denoted C1, C2, , C12, see Section 2.2). By comparing the distributions of LLM-generated takeaways vs. human-written takeaways along these comparison types, we can understand to what extent takeaways from LLM emulate those from humans.
@@ 3.3 Experiment 2: Results
Figure 5: Distributions of human, LLM zero-shot, and LLM one-shot comparison types for each layout. LLM zero-shot distributions are generally closer than one-shot distributions to human ones.
@@ 3.3.2 Alignment of Comparison Types
Table 3.3.2 further details the top three most frequent comparison types for each layout from each source while Table 3.3.2 shows precision@3. Compared to zero-shot takeaways, one-shot takeaways attained higher precision@3 for two layouts and equal precision@3 for the remaining two, indicating that the presence of example takeaways tends to allow LLMs to better capture the most afforded comparison types. While LLMs were generally good at predicting frequent comparison types across all layouts, such as C1, C11, and C12, when rarer comparison types made it to the top three, such as C3 and C4, LLMs failed to predict them. However, the frequent appearance of a less common comparison type in other layouts within the top ranks for a specific layout suggests a unique affordance by that layout for such comparisons. Failure by the LLM to capture these affordances is further evidence that LLM comparison types are generally insensitive to bar layout.
@@ 4.1 Experiment 3: Procedure, Setup, and Evaluation
To answer the second question, we generated eight new bar charts by modifying the original eight stimuli from Xiong et al. [39]. For each chart, we preserved its context while introducing a new dataset. For example, if the existing two stacked bar charts visualized “context 1 and dataset 1” and “context 2 and dataset 2”, we created two new stacked bar charts visualizing “context 1 and dataset 2” and “context 2 and dataset 1”. Next, we generated zero-shot takeaways using the optimal configuration established in Experiment 1, and the same coder from Experiment 2 coded these takeaways for comparison types. We then visualized the LLM comparison type distributions for the four example charts for each layout. To examine the effect of contexts on distributions, we computed the TVD’s between the distributions for each pair of charts showing the same dataset but different contexts and calculated the average. To examine the effect of data on distributions, we computed the TVD’s between the distributions for each pair of charts showing the same context but different datasets and calculated the average.